A company has a $25 million portfolio with a beta of 0.8. The futures price for a contract on an index is 1200. Futures contracts on $250 times the index can be traded. What trade is necessary to increase beta to 1.5?
Answer #1

Answer:

Required increase in beta = 1.5 - 0.8 = 0.7

To increase beta by 0.6 we need to go long (buy) = 0.7 * 25000000 / (1200 * $250) = 58.33 or 58 contracts.

Trade is necessary to increase beta to 1.5 is to go long (buy) 58 contracts.